# Physics-Informed Time-Integrated DeepONet: Temporal Tangent Space Operator Learning for High-Accuracy Inference

### v1

@@ 3.1 One-dimensional Heat Equation
Figure 3: Example for 1D heat equation (Mean rel. $\mathcal{L}_{2}$ error $7.60\mathrm{e}\text{-}2$) with comparison of reference data and PITI-DeepONet with explicit Euler alongside the squared difference of the computed field and the residuals during inference. Mean absolute error between squared difference and predicted residual for this example is $3.72\mathrm{e}\text{-}5$ and Pearson’s correlation coefficient is $\rho=0.9985$.
Figure 4: Total loss during training alongside mean relative $\mathcal{L}_{2}$ error on the train and validation set for full rollout (FR), autoregressive (AR), and time integration (PITI) model for the one-dimensional heat equation. The relative $\mathcal{L}_{2}$ error on both network outputs is provided for the PITI model. Hyperparameters were adapted per model as per table 2.
Figure 5: Comparison of time integration models in the vanilla form and special form for time-evolution partial differential equations including total loss and mean relative $\mathcal{L}_{2}$ error on train and validation set during training for the heat equation. Only the relative $\mathcal{L}_{2}$ error on the time derivative is provided here. The right plot shows the performance during inference with the fully trained model using an explicit euler scheme while depicting minimum, mean, and maximum value of the relative $\mathcal{L}_{2}$ error per time step.
Figure 6: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples.
Figure 7: Comparison of predictions made using time integration (PITI) with explicit euler, autoregression and full rollout for 1D heat example alongside the differences for each method.
@@ 3.2 One-dimensional Burgers’ Equation
The dataset is generated by solving Eq. 10 on a periodic spatial domain $x\in[0,1]$. The spatial domain is discretized using $s=4,096$ grid points, while outputs are captured at $n_{\text{steps}}+1=101$ evenly spaced time steps over $t\in[0,1]$. The ICs $u_{0}(x)$ are sampled from stationary Gaussian random fields (GRFs) with covariance kernel $C=\sigma^{2}(-\Delta+\tau^{2}I)^{-\gamma}$. For GRF generation, the Fourier coefficients are drawn from $\mathcal{N}(0,\sigma^{2})$ and scaled appropriately to enforce the specified covariance structure, yielding smooth, periodic initial profiles. The Burgers’ equation is advanced in time using an operator-splitting spectral method implemented with the spin function from Chebfun [26], which combines linear evolution via viscosity $\nu u_{xx}$ with nonlinear convection $uu_{x}$. The solution $u(t,x)$ is represented as a Chebyshev interpolant at each time step and stored. The corresponding time derivative $\partial_{t}u(t,x)$ is independently computed at each output time by substituting the solution into the original Burgers’ equation. Consistency checks are performed to ensure the accuracy of $\partial_{t}u(t,x)$ using both finite differences and the governing PDE. The dataset consists of $N=2,500$ realizations of $u_{0}(x)$, where each realization evolves under the Burgers’ dynamics and is evaluated on a reduced grid of $n=101$ spatial points.
@@
Figure 8: Reference and difference of predictions to reference for Burgers’ equation using time integration with fourth-order Runge-Kutta, autoregressive and full rollout.
Figure 9: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples.
Figure 10: Reference and difference of predictions to reference for Burgers’ equation using time integration with explicit Euler and different time steps.
Figure 11: Total loss during training alongside mean relative $\mathcal{L}_{2}$ error on the train and validation set for full rollout (FR), autoregressive (AR), and time integration (PITI) model for one-dimensional Burgers’ equation. The relative $\mathcal{L}_{2}$ error on both network outputs is provided for the PITI model. Hyperparameters were adapted per model as per table 2.
Figure 12: Comparison of time integration models in the vanilla form and special form for time-evolution partial differential equations alongside a purely physics-informed training including total loss and mean relative $\mathcal{L}_{2}$ error on train and validation set during training for Burgers’ equation. Only the relative $\mathcal{L}_{2}$ error on the time derivative is provided here. The right plot shows the performance during inference with the fully trained model using an explicit euler scheme while depicting minimum, mean, and maximum value of the relative $\mathcal{L}_{2}$ error per time step.
@@ 3.3 Two-dimensional Allen-Cahn Equation
Figure 13: Comparison of explict Euler as time integration (TI) methods with different time steps during inference $\Delta t$. Mean relative $\mathcal{L}_{2}$ error is shown on the left and predicted residuals on the right over all test examples.
@@
Figure 14: Reference, predictions, and absolute error in predictions with respect to reference for Allen-Cahn equation using time integration with fourth-order Runge-Kutta (RK4), autoregression and full rollout at different time points.
@@
Figure 15: Total loss during training alongside mean relative $\mathcal{L}_{2}$ error on the train and validation set for full rollout (FR), autoregressive (AR), and time integration (PITI) model for two-dimensional Allen-Cahn equation. The relative $\mathcal{L}_{2}$ error on both network outputs is provided for the PITI model. Hyperparameters were adapted per model as per table 2.
@@
Figure 16: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples.
@@ 4 Discussion
The proposed method includes a built-in mechanifor tracking residuals during field reconstruction, providing an indicative measure of prediction quality at inference time. Across all examples, the residuals exhibited a nearly perfect linear correlation with the squared difference between predictions and the reference solutions obtained using explicit Euler, highlighting the strong error-monitoring capabilities inherent in the proposed PITI-DeepONet framework. This residual tracking approach enables additional possibilities, such as coupling it with more accurate schemes when a residual threshold is exceeded or facilitating active learning by identifying whether a state lies within the training domain. It can also serve as a time-step control criterion, allowing the rejection of a step if an unlearned or unexpected state arises, ensuring consistency in the inferred behavior.
A crucial aspect of effectively employing PITI-DeepONet is ensuring that all potential states encountered during inference are adequately sampled. These states can be obtained through experiments, traditional methods such as numerical or analytical solutions, or alternative surrogate models, such as FR simulations. Incorporating expert domain knowledge can significantly enhance the efficiency and quality of the sampling process. An additional, yet unexplored, avenue is the use of multi-fidelity learning [27], where data of varying resolutions or qualities can be seamlessly included within this framework. Moreover, it has been shown that a hybrid approach, which directly incorporates observed data, can greatly improve the learning process. However, this comes with the potential requirement of time derivative data for training, which may or may not be readily available, depending on the method used to acquire the data.
It is important to emphasize that no specific assumptions were made regarding the neural network architecture, allowing for the use of flexible, problem-specific designs. This can include architectures tailored to the number of variables, such as multi-input [28] and multi-output [29], as well as multi-network setups, i.e., a single network per field variable in coupled problems [30]. In this work, DeepONets with traditional feed-forward neural networks for the branch and trunk, as well as convolutional neural networks for the branch, were utilized; however, other (physics-informed) NO architectures are equally applicable. Furthermore, this framework can be readily scaled to higher-dimensional PDEs by adopting separable architectures [31].

### v2

@@ 3.1 One-dimensional Heat Equation
Figure 3: 1D Heat equation (Mean rel. $\mathcal{L}_{2}$ error $7.60\mathrm{e}\text{-}2$): Comparison of reference data and PITI-DeepONet with explicit Euler alongside the squared difference of the computed field and the residuals during inference. Mean absolute error between squared difference and predicted residual for this example is $3.72\mathrm{e}\text{-}5$ and Pearson’s correlation coefficient is $\rho=0.9985$.
Figure 4: 1D Heat equation: Total loss during training alongside mean relative $\mathcal{L}_{2}$ error on the train and validation set for full rollout (FR), autoregressive (AR), and time integration (PITI) model. The relative $\mathcal{L}_{2}$ error on both network outputs is provided for the PITI model. Hyperparameters were adapted per model as per table 2.
Figure 5: 1D Heat equation: Comparison of time integration models in the vanilla form and special form for time-evolution partial differential equations including total loss and mean relative $\mathcal{L}_{2}$ error on train and validation set during training. Only the relative $\mathcal{L}_{2}$ error on the time derivative is provided here. The right plot shows the performance during inference with the fully trained model using an explicit Euler scheme while depicting minimum, mean, and maximum value of the relative $\mathcal{L}_{2}$ error per time step.
Figure 6: 1D Heat equation: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples.
Figure 7: 1D Heat equation: Comparison of predictions made using time integration (PITI) with explicit Euler, autoregressive and full rollout alongside the differences for each method.
@@ 3.2 One-dimensional Burgers’ Equation
The dataset is generated by solving Eq. 10 on a periodic spatial domain $x\in[0,1]$. The spatial domain is discretized using $s=4,096$ grid points, while outputs are captured at $n_{\text{steps}}+1=101$ evenly spaced time steps over $t\in[0,1]$. The ICs $u_{0}(x)$ are sampled from stationary Gaussian random fields (GRFs) with covariance kernel $C=\sigma^{2}(-\Delta+\tau^{2}I)^{-\gamma}$. For GRF generation, the Fourier coefficients are drawn from $\mathcal{N}(0,\sigma^{2})$ and scaled appropriately to enforce the specified covariance structure, yielding smooth, periodic initial profiles. The Burgers’ equation is advanced in time using an operator-splitting spectral method implemented with the spin function from Chebfun [chebfun], which combines linear evolution via viscosity $\nu u_{xx}$ with nonlinear convection $uu_{x}$. The solution $u(t,x)$ is represented as a Chebyshev interpolant at each time step and stored. The corresponding time derivative $\partial_{t}u(t,x)$ is independently computed at each output time by substituting the solution into the original Burgers’ equation. Consistency checks are performed to ensure the accuracy of $\partial_{t}u(t,x)$ using both finite differences and the governing PDE. The dataset consists of $N=2,500$ realizations of $u_{0}(x)$, where each realization evolves under the Burgers’ dynamics and is evaluated on a reduced grid of $n=101$ spatial points.
@@
Figure 8: 1D Burgers’ equation: Reference and difference of predictions to reference using time integration with fourth-order Runge-Kutta, autoregressive and full rollout.
Figure 9: 1D Burgers’ equation: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples.
Figure 10: 1D Burgers’ equation: Reference and difference of predictions to reference using time integration with explicit Euler and different time steps.
Figure 11: 1D Burgers’ equation: Total loss during training alongside mean relative $\mathcal{L}_{2}$ error on the train and validation set for full rollout (FR), autoregressive (AR), and time integration (PITI) model. The relative $\mathcal{L}_{2}$ error on both network outputs is provided for the PITI model. Hyperparameters were adapted per model as per table 2.
Figure 12: 1D Burgers’ equation: Comparison of time integration models in the vanilla form and special form for time-evolution partial differential equations alongside a purely physics-informed training including total loss and mean relative $\mathcal{L}_{2}$ error on train and validation set during training. Only the relative $\mathcal{L}_{2}$ error on the time derivative is provided here. The right plot shows the performance during inference with the fully trained model using an explicit Euler scheme while depicting minimum, mean, and maximum value of the relative $\mathcal{L}_{2}$ error per time step.
@@ 3.3 Two-dimensional Allen-Cahn Equation
Figure 13: 1D Burgers’ equation: Comparison of explicit Euler as time integration (TI) method with different time steps ($\Delta t$) during inference. Mean relative $\mathcal{L}_{2}$ error is shown on the left and predicted residuals on the right over all test examples.
@@
Figure 14: 2D Allen-Cahn equation: Reference, predictions, and absolute error in predictions with respect to reference using time integration with fourth-order Runge-Kutta (RK4), autoregressive and full rollout at different time points.
@@
Figure 15: 2D Allen-Cahn equation: Total loss during training alongside mean relative $\mathcal{L}_{2}$ error on the train and validation set for full rollout (FR), autoregressive (AR), and time integration (PITI) model. The relative $\mathcal{L}_{2}$ error on both network outputs is provided for the PITI model. Hyperparameters were adapted per model as per table 2.
@@
Figure 16: 2D Allen-Cahn equation: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples.
3.4 Ablation Studies
Table 4: Seed sensitivity analysis for the final model configurations. Each method (PITI, AR, FR) was trained with five different random seeds. For each run, the relative $\mathcal{L}_{2}$ error was computed over all test samples and summarized by its mean, minimum, and maximum at two time points corresponding to the end of the training domain and the end of the testing domain for each problem (Heat 1D: $t=0.5$ and $t=5.0$; Burgers 1D: $t=0.5$ and $t=1.0$; Allen-Cahn 2D: $t=0.5$ and $t=1.0$). Table entries report the overall mean and between-seed standard deviation across the five runs, formatted as $\mu\pm\sigma$.
Heat 1D Burgers 1D Allen-Cahn 2D
$t=0.5$ $t=5.0$ $t=0.5$ $t=1.0$ $t=0.5$ $t=1.0$
Mean
PITI $3.5\mathrm{e}\text{-}2\pm 3.9\mathrm{e}\text{-}3$ $2.2\mathrm{e}\text{-}1\pm 3.4\mathrm{e}\text{-}2$ $1.1\mathrm{e}\text{-}2\pm 6.2\mathrm{e}\text{-}4$ $1.8\mathrm{e}\text{-}2\pm 1.8\mathrm{e}\text{-}3$ $7.9\mathrm{e}\text{-}2\pm 1.4\mathrm{e}\text{-}3$ $1.2\mathrm{e}\text{-}1\pm 2.6\mathrm{e}\text{-}3$
AR $2.2\mathrm{e}\text{-}1\pm 3.3\mathrm{e}\text{-}2$ $1.5\mathrm{e}\text{-}0\pm 5.4\mathrm{e}\text{-}1$ $2.4\mathrm{e}\text{-}1\pm 8.7\mathrm{e}\text{-}2$ $6.0\mathrm{e}\text{-}1\pm 2.1\mathrm{e}\text{-}1$ $9.9\mathrm{e}\text{-}1\pm 7.2\mathrm{e}\text{-}2$ $1.1\mathrm{e}0\pm 7.0\mathrm{e}\text{-}2$
FR $5.7\mathrm{e}\text{-}3\pm 3.2\mathrm{e}\text{-}4$ $1.3\mathrm{e}0\pm 1.4\mathrm{e}\text{-}1$ $3.5\mathrm{e}\text{-}2\pm 2.0\mathrm{e}\text{-}3$ $1.2\mathrm{e}\text{-}1\pm 2.1\mathrm{e}\text{-}2$ $1.3\mathrm{e}\text{-}1\pm 2.3\mathrm{e}\text{-}2$ $2.1\mathrm{e}\text{-}1\pm 1.8\mathrm{e}\text{-}2$
Min
PITI $3.8\mathrm{e}\text{-}3\pm 4.2\mathrm{e}\text{-}4$ $2.8\mathrm{e}\text{-}2\pm 1.7\mathrm{e}\text{-}3$ $2.5\mathrm{e}\text{-}3\pm 3.9\mathrm{e}\text{-}4$ $4.4\mathrm{e}\text{-}3\pm 5.9\mathrm{e}\text{-}4$ $5.3\mathrm{e}\text{-}2\pm 1.5\mathrm{e}\text{-}3$ $7.7\mathrm{e}\text{-}2\pm 3.6\mathrm{e}\text{-}3$
AR $3.2\mathrm{e}\text{-}2\pm 8.3\mathrm{e}\text{-}3$ $1.3\mathrm{e}\text{-}1\pm 5.9\mathrm{e}\text{-}2$ $4.1\mathrm{e}\text{-}2\pm 7.0\mathrm{e}\text{-}3\mathdollar$ $1.0\mathrm{e}\text{-}1\pm 3.3\mathrm{e}\text{-}2$ $1.8\mathrm{e}\text{-}1\pm 3.6\mathrm{e}\text{-}2$ $1.9\mathrm{e}\text{-}1\pm 4.3\mathrm{e}\text{-}2$
FR $1.4\mathrm{e}\text{-}3\pm 3.4\mathrm{e}\text{-}4$ $6.3\mathrm{e}\text{-}1\pm 1.3\mathrm{e}\text{-}1$ $7.0\mathrm{e}\text{-}3\pm 7.2\mathrm{e}\text{-}4$ $3.9\mathrm{e}\text{-}2\pm 1.0\mathrm{e}\text{-}2$ $4.0\mathrm{e}\text{-}2\pm 7.1\mathrm{e}\text{-}3$ $7.1\mathrm{e}\text{-}2\pm 1.3\mathrm{e}\text{-}2$
Max
PITI $1.7\mathrm{e}\text{-}1\pm 3.4\mathrm{e}\text{-}2$ $2.5\mathrm{e}\text{-}0\pm 6.6\mathrm{e}\text{-}1$ $1.9\mathrm{e}\text{-}1\pm 3.7\mathrm{e}\text{-}2$ $2.5\mathrm{e}\text{-}1\pm 7.1\mathrm{e}\text{-}2$ $1.1\mathrm{e}\text{-}1\pm 6.0\mathrm{e}\text{-}3$ $1.8\mathrm{e}\text{-}1\pm 1.2\mathrm{e}\text{-}2$
AR $9.7\mathrm{e}\text{-}1\pm 2.1\mathrm{e}\text{-}1$ $3.2\mathrm{e}1\pm 1.7\mathrm{e}1$ $1.8\mathrm{e}0\pm 1.0\mathrm{e}0$ $5.2\mathrm{e}0\pm 3.5\mathrm{e}0$ $3.3\mathrm{e}0\pm 4.1\mathrm{e}\text{-}1$ $2.9\mathrm{e}0\pm 3.8\mathrm{e}\text{-}1$
FR $2.9\mathrm{e}\text{-}2\pm 5.9\mathrm{e}\text{-}3$ $2.1\mathrm{e}1\pm 2.5\mathrm{e}0$ $2.5\mathrm{e}\text{-}1\pm 3.7\mathrm{e}\text{-}2$ $5.3\mathrm{e}\text{-}1\pm 9.3\mathrm{e}\text{-}2$ $3.4\mathrm{e}\text{-}1\pm 6.8\mathrm{e}\text{-}2$ $4.7\mathrm{e}\text{-}1\pm 6.6\mathrm{e}\text{-}2$
In addition to the investigations described above and the network architectures explored during hyperparameter optimization, the final configurations were trained using five different random seeds. For evaluation, the relative $\mathcal{L}_{2}$ error was computed for each test sample, and its mean, minimum, and maximum were recorded at two representative time points: (i) the end of the training time horizon and (ii) the end of the test time horizon. Across the five seeded trainings, we report the grand mean and the between-seed standard deviation (formatted as $\mu\pm\sigma$) of these summary statistics in Table 4.
Overall, the between-seed standard deviations are typically at least one order of magnitude smaller than the corresponding error levels. In the remaining cases, the variability is either noticeably smaller while remaining within the same order of magnitude, or the absolute error is already sufficiently large that seed-induced variation is negligible in comparison. This indicates that the reported performance is robust with respect to the choice of random seed and that seed effects do not materially affect the conclusions drawn in the previous sections.
@@ 4 Discussion
The proposed method includes a built-in mechanism for tracking residuals during field reconstruction, providing an indicative measure of prediction quality at inference time. Across all examples, the residuals exhibited a nearly perfect linear correlation with the squared difference between predictions and the reference solutions obtained using explicit Euler, highlighting the strong error-monitoring capabilities inherent in the proposed PITI-DeepONet framework. This residual tracking approach enables additional possibilities, such as coupling it with more accurate schemes when a residual threshold is exceeded or facilitating active learning by identifying whether a state lies within the training domain. It can also serve as a time-step control criterion, allowing the rejection of a step if an unlearned or unexpected state arises, ensuring consistency in the inferred behavior.
A crucial aspect of effectively employing PITI-DeepONet is ensuring that all potential states encountered during inference are adequately sampled. These states can be obtained through experiments, traditional methods such as numerical or analytical solutions, or alternative surrogate models, such as FR simulations. Incorporating expert domain knowledge can significantly enhance the efficiency and quality of the sampling process. An additional, yet unexplored, avenue is the use of multi-fidelity learning [howard2023multifidelity], where data of varying resolutions or qualities can be seamlessly included within this framework. Moreover, it has been shown that a hybrid approach, which directly incorporates observed data, can greatly improve the learning process. However, this comes with the potential requirement of time derivative data for training, which may or may not be readily available, depending on the method used to acquire the data.
It is important to emphasize that no specific assumptions were made regarding the neural network architecture, allowing for the use of flexible, problem-specific designs. This can include architectures tailored to the number of variables, such as multi-input [jin2022mionet] and multi-output [Lu2022], as well as multi-network setups, i.e., a single network per field variable in coupled problems [Cai2021]. In this work, DeepONets with traditional feed-forward neural networks for the branch and trunk, as well as convolutional neural networks for the branch, were utilized; however, other (physics-informed) NO architectures are equally applicable. Furthermore, this framework can be readily scaled to higher-dimensional PDEs by adopting separable architectures [Mandl2025].